How much should be invested at 8% compounded semiannually in order to have $5000 at the end of 8 years.: How much should be invested at 8% compounded semiannually in order to have $5000 at the end of 8 years.
Answer by checkley77(1586)

X(1+.08)^8=5000
X(1.08)^8=5000
1.85X=5000
X=5000/1.85
X=$2,702.70 IS THE AMOUNT INVESTED.